### v2

# A fast Bellman algorithm for the real Monge–Ampère equation

Abstract.
In this paper, we introduce a new numerical algorithm for solving the Dirichlet problem for the real Monge–Ampère equation. The idea is to represent the non-linear Monge–Ampère operator as an infimum of a class of linear elliptic operators and use Bellman’s principle to construct a numeric scheme for approximating the operator attaining this infimum.
Moreover, we prove convergence of the proposed algorithm (under suitable technical assumptions) and discuss its strengths and weaknesses. We also demonstrate the performance of the method on several examples with various degrees of regularity and degeneracy and compare the results to two existing methods. Our method runs considerably faster than the ones used for comparison, improving the running time by a factor of 3–10 for smooth, strictly convex examples, and by a factor of 20–100 or more for mildly degenerate examples.
Key words and phrases:
non-linear elliptic differential equation, real Monge–Ampère equation, linearized Monge–Ampère equation, Bellman principle, fixed point algorithm
2020 Mathematics Subject Classification:
Primary 65N06; Secondary 35J60, 35J96, 65N12
1. Introduction
The Monge–Ampère equation is a fully non-linear partial differential equation, originating from the study of surfaces with prescribed curvature which goes back to Monge in the late 18th and Ampère in the early 19th century. Since then, variations of this equation have found a number of applications, for example in optimal transport [13, 28] and Riemannian geometry (the local isometric embedding problem [25]). The equation is also of great theoretical interest and is often used as a model example of a fully non-linear PDE. The complex version of the Monge–Ampère equation plays a central role in pluripotential theory, i.e. in the study of plurisubharmonic functions on domains in $\mathbb{C}^{n}$ and more recently on Kähler manifolds [17].
In this paper, we introduce a novel numerical method for solving the Dirichlet problem for the real Monge–Ampère equation, based on Bellman’s principle: representing the non-linear operator as an infimum of elliptic linear partial differential operators. We construct a numeric iterative scheme for approximating the operator attaining this infimum.
Under suitable technical assumptions, as well as the assumption that our algorithm actually produces a sequence of strongly convex approximants, we prove convergence, see Theorem 4.1 for details. Also, any convex approximants always majorises the solution to the original Monge–Ampère equation.
Numerical experiments illustrate that the method works well in settings where the solution is smooth and strictly convex. In these cases, it outperforms recently published methods by a factor of 3–10 in execution time. For mildly singular equations, the speed-up is even better: in some cases improving the running time by a factor of 100 or more. In the fully degenerate case, when the right hand side of the equation vanishes on a non-empty open set, our proposed method does not converge in general. We intend to return to the fully degenerate problem in a subsequent paper.
2. Background
The Dirichlet problem for the Monge–Ampère operator can be formulated as
(2.1) $\begin{cases}\det(D^{2}u(x))=f(x),&x\in\Omega,\\ u(x)=\phi(x),&x\in\partial\Omega,\end{cases}$
where $\Omega\subset\mathbb{R}^{n}$ is a convex domain, $f\geq 0$ and the boundary values of $u$ are suitably interpreted. Here, $D^{2}u=(u_{jk})$ denotes the Hessian of $u$. By restricting our attention to convex solutions of (2.1), the equation is (degenerate) elliptic, and we will only consider such solutions. If we impose some extra conditions on $\Omega$, $f$ and $\phi$, we can guarantee that the equation has a unique solution. Note that the left-hand side, the Monge–Ampère operator, $\operatorname{MA}(u):=\det(D^{2}u(x))$ makes perfect sense for $C^{2}$ functions, and it can be interpreted pointwise also for functions in $W_{\text{loc}}^{2,n}$. For functions of lower regularity, the Monge–Ampère operator must be viewed in a generalized sense, and since the operator is fully non-linear, there is no direct interpretation in the sense of distributions. However, convex functions are automatically locally Lipschitz and it turns out that we can define the Monge–Ampère measure (in the sense of Aleksandrov) of any convex function $u$, by
$\mu_{u}(E)=\operatorname{Leb}(\partial u(E))$
for any Borel subset $E\subset\Omega$, where $\operatorname{Leb}$ denotes the Lebesgue measure, $\partial u(E)=\cup_{x\in E}\partial u(x)$, and $\partial u(x)$ is the set of all slopes of the supporting hyperplanes to the graph of $u$ at $x$. If $u\in C^{2}$ and $D^{2}u$ is positive semi-definite, then $\mu_{u}(E)=\det(D^{2}u)$. We should mention, that in contrast to the real setting, it is impossible to define the complex Monge–Ampère operator for every plurisubharmonic function. We refer to the monograph by Guedj and Zeriahi [17] for the theory of complex Monge–Ampère equations.
There is a vast number of analytic results in the literature concerning existence and regularity of solutions to (2.1). A complete overview is outside the scope of this paper, but for example, it follows from results by Aleksandrov [1] and Bakelman [3] from the 1950’s that if $\Omega$ is assumed to be a bounded strictly convex domain, $f$ is a finite Borel measure on $\Omega$, and $\phi\in C(\partial\Omega)$, then (2.1) has a unique convex solution $u$ (in the sense of Aleksandrov) with $u\in C(\bar{\Omega})$.
Furthermore, if $\Omega$ is strictly convex (or uniformly convex in the unbounded setting) with $C^{k+2,\alpha}$ boundary, $0<\phi\in C^{k+2,\alpha}(\partial\Omega)$ and $f\in C^{k,\alpha}(\bar{\Omega})$ for some integer $k\geq 2$ and some $10<\alpha<$ 1, then the Dirichlet problem (2.1) has a unique convex solution $u\in C^{k+2,\alpha}(\bar{\Omega})$, see [9, 21, 22].
If we merely assume that $\Omega$ is convex, not necessarily strictly convex, the situation is a little less straight-forward, and we need to assume that $\phi$ is convex on all line segments in $\partial\Omega$ to guarantee existence of solutions to (2.1). Similarly, if $f$ is only assumed to be non-negative instead of strictly positive, the solutions to the Dirichlet problem will generally be less smooth, even in strictly convex domains. If $f$ has zeros, the equation is called degenerate. See for example Błocki [7] for results in this direction, where $C^{1,1}$-regularity of the solution is shown under suitable technical assumptions. We will come back to this phenomenon in example 6.2.3.
For an accessible and thorough overview of the Aleksandrov solutions as well as results on classical solutions including the ones mentioned, we refer to the recent monographs by Figalli [14] and Le [24].
There have been a number of suggested numerical methods for solving the Dirichlet problem for the Monge–Ampère operator. As far as we are aware, the first approach was due to Oliker and Prussner (1988) [27] who proposed a direct discretization of the Aleksandrov interpretation of the Monge–Ampère operator. Since then, a number of other approaches to solving Monge–Ampère equations have been tried: finite difference discretization, often using wide stencils to obtain stable and consistent methods [5, 6, 11, 12, 26], variational methods [4, 19], fixpoint methods [5, 18] as well as finite elements [20, 23] only to mention a few.
3. Bellman’s principle
Our method relies on the following linear algebra result, sometimes called Bellman’s principle. In the setting of complex Monge–Ampère equations, this idea was introduced by Gaveau [15] and has later found several theoretical applications. We will formulate the result in the real setting and for convenience, we provide a proof of Bellman’s principle, similar to the approach taken by Wiklund [29].
Theorem 3.1.
Let $M$ be a positive semi-definite symmetric $n\times n$ real matrix. Then
(3.1) $(\det M)^{1/n}=\frac{1}{n}\inf_{B\in\mathcal{B}}\operatorname{tr}(BM),$
where $\operatorname{tr}(\cdot)$ denotes the trace of a matrix, and $\mathcal{B}$ denotes the set of positive definite symmetric $n\times n$-matrices with determinant $1$. Also, if $M$ is positive definite, the infimum is attained by some matrix $B$.
Proof.
First, we show that
$(\det M)^{1/n}\leq\frac{1}{n}\inf_{B\in\mathcal{B}}\operatorname{tr}(BM).$
To this end we observe that for any symmetric and positive definite matrix $B$, we have
$\begin{split}\det(BM-\lambda I)&=\det(B^{1/2}B^{1/2}MB^{1/2}B^{-1/2}-\lambda I)\\ &=\det(B^{1/2})\det(B^{1/2}MB^{1/2}-\lambda I)\det(B^{-1/2})\\ &=\det(B^{1/2}MB^{1/2}-\lambda I).\end{split}$
Since $B^{1/2}MB^{1/2}$ is symmetric and positive semi-definite, it can be diagonalised by a unitary matrix $Q$ and the eigenvalues of $BM$, $\{\lambda_{k}\}_{1\leq k\leq n}$, are non-negative. Moreover, for $B$ Hermitian with $\det(B)=1$ we have
$\begin{split}\det(M)=\det(BM)&=\prod_{1\leq k\leq n}\lambda_{k}\leq\Big{(}\frac{1}{n}\sum_{1\leq k\leq n}\lambda_{k}\Big{)}^{n}\\ &=\Big{(}\frac{1}{n}\operatorname{tr}(Q^{*}B^{1/2}MB^{1/2}Q)\Big{)}^{n}\\ &=\Big{(}\frac{1}{n}\operatorname{tr}(B^{1/2}QQ^{*}B^{1/2}M)\Big{)}^{n}=\Big{(}\frac{1}{n}\operatorname{tr}(BM)\Big{)}^{n},\end{split}$
using the inequality between the arithmetic and geometric mean. Next, we show that the infimum is attained for positive definite matrices. If $M$ is a positive definite matrix then $B\coloneqq(\det M)^{1/n}M^{-1}$ is a well-defined, positive definite matrix with $\det B=\det M\cdot\det M^{-1}=1$ which satisfies the statement. Indeed,
$\operatorname{tr}{BM}=\operatorname{tr}{((\det M)^{1/n}M^{-1}M)}=(\det M)^{1/n}n.$
If $M$ is diagonal and positive semi-definite, let $p$ be the rank of $M$. Clearly, $p<n$. Without loss of generality, we can assume that $\lambda_{k}=0$ for $k\in\{p+1,\dots,n\}$. In this case, the equality can be obtained by choosing any constant $c>0$ and constructing a diagonal matrix $B$ with $b_{kk}\coloneqq c$ for $k\in\{1,\dots,p\}$ and $b_{kk}\coloneqq c^{-p/(n-p)}$ for $k\in\{p+1,\dots n\}$ so that $\det(B)=1$. Then, we have
$\begin{split}\operatorname{tr}(BM)&=\sum_{k=1}^{p}b_{kk}\lambda_{k}=c\cdot\operatorname{tr}(M)\implies\\ &\qquad\inf_{B\in\mathcal{B}}\operatorname{tr}{(BM)}\leq\inf_{c>0}c\cdot\operatorname{tr}(M)=0=n\det(M)^{1/n}.\end{split}$
Combining this with $n\det(M)^{1/n}\leq\inf_{B\in\mathcal{B}}\operatorname{tr}(BM)$ derived at the beginning, we obtain Equation (3.1). For a positive semi-definite matrix $M$ that is not diagonal, it suffices to notice that there exists $T$ such that $M=T^{*}\Lambda T$, where $\Lambda$ is a diagonal matrix with the eigenvalues of $M$ on the diagonal. By the above, we have
$\begin{split}(\det M)^{1/n}&=(\det\Lambda)^{1/n}=\frac{1}{n}\inf_{B\in\mathcal{B}}\operatorname{tr}(B\Lambda)=\frac{1}{n}\inf_{B\in\mathcal{B}}\operatorname{tr}(BTMT^{*})\\ &=\frac{1}{n}\inf_{B\in\mathcal{B}}\operatorname{tr}(T^{*}BTM)=\frac{1}{n}\inf_{B^{\prime}\in\mathcal{B}}\operatorname{tr}(B^{\prime}M).\qed\end{split}$
Using Bellman’s principle, we can describe the Monge–Ampère operator (of a sufficiently smooth function) as an infimum over a class of elliptic partial differential operators. More precisely:
Corollary 3.2.
Let $\Omega$ be a domain in $\mathbb{R}^{n}$ and assume that $u\in C^{2}(\Omega)$ is convex. Then
$(\operatorname{MA}(u)(x))^{1/n}=\frac{1}{n}\inf_{B\in\mathcal{B}}\operatorname{tr}\big{(}B\,D^{2}u(x))\big{)}.$
In particular, if $u$ is strongly convex, the infimum is attained by the $\mathcal{B}$-valued function
$B(x)=(\det D^{2}u)^{1/n}(D^{2}u)^{-1}(x).$
Note that, for every choice of $\mathcal{B}$-valued function $B$, the operator
$u\mapsto\operatorname{tr}\big{(}B(x)D^{2}u(x)\big{)}$
is a linear elliptic partial differential operator, and if we somehow know the associated $B(x)$ corresponding to the solution $u$ of the Dirichlet problem (2.1), then we can recover $u$ by solving the linear elliptic Dirichlet problem
(3.2) $\begin{cases}\operatorname{tr}(B(x)\,D^{2}u(x))=n\,\big{(}f(x)\big{)}^{1/n},&x\in\Omega,\\ u(x)=\phi(x),&x\in\partial\Omega,\end{cases}$
for which there are highly efficient numerical algorithms. Let us mention that with this approach we have the following monotonicity result.
Proposition 3.3.
Let $\Omega\subset\mathbb{R}^{n}$ be a convex domain with $C^{2,\alpha}$ boundary, $\phi\in C^{2,\alpha}(\partial\Omega)$, $f^{1/n}\in C^{0,\alpha}(\Omega)$ and let $B\in C^{0,\alpha}(\Omega)$ be a $\mathcal{B}$-valued function on $\Omega$. Assume that the solution to the above Dirichlet problem (Eq. (3.2)) is convex. Then $u\geq u_{\text{MA}}$ where $u_{\text{MA}}$ denotes the solution to the corresponding Monge–Ampère equation (2.1).
Proof.
By standard regularity theory for elliptic PDEs, $u\in C^{2,\alpha}$, so by Corollary 3.2, $\operatorname{MA}(u)\leq f$. Hence, by the comparison principle for the real Monge–Ampère operator (see for example [24, Theorem 3.21]), it follows that $u\geq u_{\text{MA}}$. ∎
Unfortunately, it seems like there are no known effective characterizations of when the solution to Equation (3.2) is convex, even in the case when $B\equiv I$, i.e. for Poisson’s equation. Our numerical experiments suggest that our proposed algorithm (as described in Section 5) produces a decreasing sequence of approximants converging to the solution of (2.1), at least once one of the approximants is convex. So far, we have not been able to prove this in full generality, but under suitable technical conditions and the assumption that the method produces a sequence of strictly convex approximants, we can show convergence, see Theorem 4.1.
We should point out that there is no proof of convergence for most methods that have been proposed for the Monge–Ampère equation. In particular, this is the case for the M1 method we compare our method to in Section 6. The convergence (under suitable conditions) of method M2, also discussed in Section 6, has been recently proved in [8, Theorem 4.1], at least in the case when the boundary data is equal to zero. Notable methods with a proof of convergence are the direct discretization of the Aleksandrov construction proposed in [27] and the wide stencil finite difference discretizations [26] which are also known to converge (to a viscosity solution).
In the remainder of the paper, we will restrict our attention to the two-dimensional case, $n=2$, but we expect the method to be competitive also in higher dimensions.
4. Proof of convergence
The constructive proof of Bellman’s principle, i.e. Theorem 3.1, suggests a method of constructing a sequence of linear Dirichlet problems of the form  (3.2) converging to the Dirichlet problem for the Monge–Ampère equation (2.1). The result is the theorem below.
Theorem 4.1.
Let $\alpha\in(0,1)$ and let $\Omega\subset\mathbb{R}^{2}$ be a strictly convex, bounded, $C^{2,\alpha}$ domain, $\phi\in C^{2,\alpha}(\bar{\Omega})$ and $f\in C^{0,\alpha}(\bar{\Omega})$ where $f>0$ on $\bar{\Omega}$. Assume that there exist strongly convex solutions $\{u_{k}\}_{k\geq 0}\subset C^{2,\alpha}(\bar{\Omega})$ to the following linear Dirichlet problems
(4.1) $\begin{cases}\operatorname{tr}(B_{k}(x)D^{2}u_{k}(x))=2\sqrt{f(x)},&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega,\end{cases}$
where $B_{k}=\sqrt{\det D^{2}u_{k}}\,(D^{2}u_{k})^{-1}$ for $k\geq 1$ and $B_{0}$ is a $\mathcal{B}$-valued function with coefficients $(B_{0})_{ij}\in C^{0,\alpha}(\bar{\Omega})$. Moreover, denote by $\lambda_{k},\Lambda_{k}$ the eigenvalues of $D^{2}u_{k}$ and assume that there exists $\epsilon>0$ such that $\epsilon\leq\lambda_{k},\Lambda_{k}$ for all $k$. Then, there exists a subsequence of the solutions $\{u_{k}\}_{k\geq 0}$ converging uniformly to a convex function $\tilde{u}\in C^{1,1}(\bar{\Omega})$ solving the Monge–Ampère equation (2.1).
Before proving the above theorem, let us first discuss the requirement of the existence of solutions to (4.1) and their strong convexity. Under the regularity assumptions stated above on the domain $\Omega$, the functions $f$ and $\phi$, and the coefficients $(B_{0})_{ij}$, standard theory of linear elliptic PDEs (see Gilbarg and Trudinger [16, Theorem 6.14]) provides existence and uniqueness of a solution $u_{0}\in C^{2,\alpha}(\bar{\Omega})$.
Now, let us investigate solvability and well-definiteness of the next Dirichlet problem in the sequence. We would like to generalise this argument for any $k\geq 0$. To this end, assume the existence of the solution $u_{k}\in C^{2,\alpha}(\bar{\Omega})$.
Observe that $B_{k+1}\coloneqq\sqrt{\det D^{2}u_{k}}(D^{2}u_{k})^{-1}$ is not necessarily well-defined, nor is the next differential operator $L_{k+1}(v)\coloneqq\operatorname{tr}(B_{k+1}D^{2}v)$ elliptic. Consequently, the solution to the corresponding Dirichlet problem is not guaranteed.
However, strong convexity of $u_{k}$ renders the next differential operator $L_{k+1}$ well-defined and elliptic. To this end, note that $D^{2}u_{k}\in C^{0,\alpha}(\bar{\Omega})$ and thus $\det(D^{2}u_{k})\in C^{0,\alpha}(\bar{\Omega})$. Hence, both eigenvalues are bounded from above. Now, if $\epsilon\leq\lambda_{k},\Lambda_{k}$ it follows that $D^{2}u_{k}$ is positive definite and hence invertible with a positive determinant. Thus $B_{k+1}$ is well-defined and by construction it is also positive definite. Furthermore, the regularity of the solution, i.e. $u_{k}\in C^{2,\alpha}(\bar{\Omega})$ results in $(B_{k})_{ij}\in C^{0,\alpha}(\bar{\Omega})$. Thus $L_{k+1}$ is elliptic with the necessary regularity of coefficients for the corresponding Dirichlet problem to have a unique solution $u_{k+1}\in C^{2,\alpha}(\bar{\Omega})$ (by another application of Theorem 6.14 in Gilbarg and Trudinger).
Proof.
We begin with defining $\nu_{k}\coloneqq\det D^{2}u_{k}(x)$ and noting that $\{u_{k}\}_{k\geq 0}$ are solutions to the following non-linear Dirichlet problems:
(4.2) $\begin{cases}\det D^{2}u_{k}(x)=\nu_{k}(x),&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega,\\ u_{k}\text{ strongly convex.}\end{cases}$
Moreover, the existence of a uniform lower bound of the eigenvalues $\lambda_{k},\Lambda_{k}$ implies the existence of a uniform upper bound.
Indeed, recall that since $B_{k+1}=\sqrt{\det D^{2}u_{k}}\,(D^{2}u_{k})^{-1}$, it is positive definite with $\det B_{k+1}=1$, for every $k\geq 0$, and we can use the Corollary to Bellman’s principle (Corollary 3.2) to conclude that
(4.3) $2\epsilon\leq 2\sqrt{\det D^{2}u_{k}}=\operatorname{tr}{B_{k+1}D^{2}u_{k}}\leq\operatorname{tr}{B_{k}D^{2}u_{k}}=2\sqrt{f}\text{ on }\bar{\Omega}.$
Thus,
(4.4) $\epsilon\leq\lambda_{k},\Lambda_{k}\leq\|\sqrt{f}\|_{C(\bar{\Omega})}/\epsilon,\qquad\text{for $k\geq 0$.}$
Consequently, $\{\nu_{k}\}_{k}$ can be interpreted as finite Borel measures with $\nu_{k}(\Omega)\leq(\|f\|_{C(\Omega)}/\epsilon)^{2}$ and $\{u_{k}\}_{k}$ as Aleksandrov solutions to (4.2). Moreover, there exists a finite Borel measure $\nu$ and a subsequence $\{\nu_{k}\}_{k\geq 0}$ (after re-enumeration) such that $\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}\nu_{k}$. By compactness of Aleksandrov solutions to the Dirichlet problem (4.2) ([24, Theorem 3.35]), we can conclude that there exists $\tilde{u}\in C(\bar{\Omega})$ and a subsequence $\{u_{k}\}_{k\geq 0}$ (after re-enumeration) such that $u_{k}\to\tilde{u}$ uniformly and $\tilde{u}$ satisfies the following
$\begin{cases}\det D^{2}\tilde{u}(x)=\nu(x),&x\in\Omega,\\ \tilde{u}(x)=\phi(x),&x\in\partial\Omega,\\ \tilde{u}\text{ convex.}\end{cases}$
It remains to show that $\tilde{u}$ solves the initial Monge–Ampère equation (2.1) by showing that $f=\nu$ as measures. To this end, we observe that all partial derivatives $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ are uniformly bounded from above and below.
Indeed, we can use Rayleigh quotients to conclude that $\{(D^{2}u_{k})_{ii}\}_{k\geq 0}$ are bounded by the eigenvalues. The existence of uniform lower and upper bounds follows from Equation (4.4). Now, using $\det D^{2}u_{k}=(D^{2}u_{k})_{ii}(D^{2}u_{k})_{jj}-(D^{2}u_{k})^{2}_{ij}$ and Equation (4.3), we see that $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ are also uniformly bounded. Thus we can choose a subsequence $\{u_{k}\}_{k\geq 0}$ (after re-enumeration) such that $u_{k}\to\tilde{u}$ uniformly and $\{(D^{2}u_{k})_{ii}\}_{k\geq 0}$, $\{(D^{2}u_{k})_{jj}\}_{k\geq 0}$ and $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ have a weak limit.
To show that $f=\nu$ in measures, we recall Equation (4.1) and $\operatorname{tr}(B_{k+1}D^{2}u_{k})=2\sqrt{\det(D^{2}u_{k})}\coloneqq 2\sqrt{\nu_{k}}$ and establish the following
$2\sqrt{f}$ $=\operatorname{tr}(B_{k}D^{2}u_{k})=\operatorname{tr}(B_{k}-B_{k+1})D^{2}u_{k}+\operatorname{tr}(B_{k+1}D^{2}u_{k})$
$=\underbrace{\operatorname{tr}(B_{k}-B_{k+1})D^{2}u_{k}}_{(*)}+\underbrace{2\sqrt{\nu_{k}}}_{(**)}.$
Furthermore, recall that $B_{k}=\sqrt{\det D^{2}u_{k-1}}(D^{2}u_{k-1})^{-1}$, so Cramer’s rule gives
$\begin{split}(B_{k})_{11}&=(D^{2}u_{k-1})_{22}/\sqrt{\det D^{2}u_{k-1}},\\ (B_{k})_{22}&=(D^{2}u_{k-1})_{11}/\sqrt{\det D^{2}u_{k-1}},\\ (B_{k})_{12}&=-(D^{2}u_{k-1})_{21}/\sqrt{\det D^{2}u_{k-1}}.\end{split}$
Consequently, $(*)\xrightarrow{w^{*}}0$ as $k\to\infty$ since the denominators are uniformly bounded away from 00 by (4.3). Furthermore ()(**)( ∗ ∗ ) is also uniformly bounded by Equation (4.3). This yields
$f-\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(f-\nu_{k})=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(\sqrt{f}-\sqrt{\nu_{k}})(\sqrt{f}+\sqrt{\nu_{k}})=0.$
Since $u_{k}$ converge uniformly and all the second-order derivatives are uniformly bounded, it follows from the Arzelà–Ascoli theorem that $\tilde{u}\in C^{1,1}(\bar{\Omega})$. ∎
5. Description of our algorithm
Theorem 4.1 provides the idea for our numerical method, which we name Bellman’s algorithm and describe below.
First, we discretize the domain $\Omega$ using a rectangular $N\times N$ mesh. For simplicity of the implementation, we only consider the case where $\Omega$ is a bounded rectangle in $\mathbb{R}^{2}$, despite the fact that these domains are not strictly convex, and the theoretical framework for solving Monge–Ampère equations on rectangles is less well-developed.
Next, we construct a sequence of functions $u_{k}$ where $u_{k}$ solves the discretized version of the following Dirichlet problem, using a narrow stencil second order finite difference solver:
(5.1) $\begin{cases}\operatorname{tr}(B_{k}(x)D^{2}u_{k}(x))=2\sqrt{f(x)},&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega.\end{cases}$
Here $B_{k}$ is a sequence of $\mathcal{B}$-valued functions constructed as follows:
  1. (1)
    $u_{0}$ is obtained by choosing $B_{0}=I$, i.e. the identity matrix. We observe that this renders equation (5.1) a Dirichlet problem for the Poisson equation.
  2. (2)
    for $k\geq 1$ we loop over the mesh and set $B_{k}\coloneqq\sqrt{\det D^{2}u_{k-1}}(D^{2}u_{k-1})^{-1}.$ If the computed $D^{2}u_{k-1}$ is not positive definite, we mark the corresponding point $x$, and temporarily take $B_{k}(x)$ as the identity matrix.
  3. (3)
    We loop over the mesh a second time and replace $B_{k}(x)$ at the marked points with the (normalized to determinant 1) convex combination of the computed Bellman matrices corresponding to the nearest grid points in horizontal and vertical direction for which $D^{2}u_{k-1}$ was positive definite.
We terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or when the number of iterations exceeds 10 0001000010\,00010 000.
If the solution of the original Monge–Ampère equation is strictly convex, we expect the constructed functions $u_{k}$ to be strictly convex as well, at least when $k$ sufficiently large, but it can happen that convexity fails in the first few steps – see Example 6.1.2 – even when the final solution is strictly convex. The interpolation step (3) is necessary to assure convergence and in practice it pushes the approximants $u_{k}$ to become more and more convex as $k$ increases. We will illustrate this phenomenon more clearly in Section 6.
If the solution of the original Monge–Ampère equation is not strictly convex, i.e. if $f$ has zeros, we cannot expect our method to converge. In practice, our proposed algorithm still performs very well when the set of degeneracy is small, for example if $f$ has an isolated zero, or even if the zero-set of $f$ is one-dimensional (cf. Example 6.2.2).
In the fully degenerate case, where $f$ vanishes on a non-empty open set or even vanishes identically, our proposed method fails to converge, or at least converges very slowly (cf. Example 6.3.1). We intend to return to the degenerate Monge–Ampère equations in a subsequent paper.
In our implementation of the Bellman method, we use the Python package FinDiff [2] to perform the necessary computations. Also, for efficiency reasons, the computations in step (2) are performed in parallel, utilizing NumPy’s tensor operations, thus removing the need for an explicit loop over the mesh elements.
For comparison, we also implemented two methods (M1 and M2) proposed by Benamou et al. [5].
The M2 method is a fixed-point method which in spirit is somewhat similar to our method. The idea is to solve a sequence of Dirichlet problems for the Poisson equation, and update the right hand side in each iteration, whereas we instead change the partial differential operator at each step, keeping the right hand side constant. For detailed descriptions of these methods, see [5], but in short, for M2 we solve
(5.2) $\begin{cases}\Delta u_{k}=g_{k}(x),&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega.\end{cases}$
where $g_{0}=2\sqrt{f}$ and $g_{k+1}=T(g_{k})=\sqrt{(\Delta u_{k})^{2}+2(f-\det(D^{2}u_{k}))}$.
It’s not too hard to show that the solution to the original Monge–Ampère equation (2.1) is obtained when $g$ is a fixed point to the operator  $T$. Note that our proposed algorithm can also be viewed as a fixed-point method: If $u$ is a strictly convex solution of the Monge–Ampère equation (2.1), and $B(x)$ is the associated Bellman function, the solution of (5.1) reproduces $u$.
Method M1 is different. Using a central difference discretization of the second order derivatives of $u$, the discretized version of the Monge–Ampère equation
$(\mathcal{D}^{2}_{xx}u_{ij})(\mathcal{D}^{2}_{yy}u_{ij})-(\mathcal{D}^{2}_{xy}u_{ij})^{2}=f_{ij}$
is rewritten as a quadratic equation for $u_{ij}$ as
(5.3) $4(a_{1}-u_{ij})(a_{2}-u_{ij})-\frac{1}{4}(a_{3}-a_{4})^{2}=h^{4}f_{ij},$
where
$\begin{cases}a_{1}=\frac{1}{2}(u_{i+1,j}+u_{i-1,j}),\\ a_{2}=\frac{1}{2}(u_{i,j+1}+u_{i,j-1}),\\ a_{3}=\frac{1}{2}(u_{i+1,j+1}+u_{i-1,j-1}),\\ a_{4}=\frac{1}{2}(u_{i-1,j+1}+u_{i+1,j-1}).\end{cases}$
By solving (5.3) for $u_{ij}$, selecting the smaller solution of the quadratic equation to impose local convexity, we have an iterative scheme for M1:
(5.4) $u_{ij}=\frac{1}{2}\left(a_{1}+a_{2}+\sqrt{(a_{1}-a_{2})^{2}+\tfrac{1}{4}(a_{3}-a_{4})^{2}+h^{4}f_{ij}}\,\right),$
which is then solved using Gauss–Seidel iteration, keeping the boundary values fixed and updating one $u_{ij}$ at the time, using (5.4). We should note that this is computationally expensive, especially in our NumPy implementation where explicit loops are slow, so M1, and to some extent our proposed method, are the ones that would benefit the most from an optimized implementation.
In our comparison we chose the same termination criterion for all methods, i.e. to terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or the number of iterations exceeds 10 0001000010\,00010 000 (for M1, we allow up to 300 000300000300\,000300 000 iterations, where one iteration means updating each $u_{ij}$ once).
6. Numerical experiments
In this section, we show the results of testing the Bellman algorithm on several examples with various degrees of smoothness and regularity of the right-hand side of the Monge–Ampère equation, $f$, on the square $\Omega=[-1,1]^{2}$, unless otherwise stated. Moreover, we reproduce the algorithms M1 and M2 presented in [5].
Then we compare the performance of both algorithms in terms of the approximation of the exact solution (restricted to the mesh) $\tilde{u}$ with the numerical solution $u$, i.e. $\|u-\tilde{u}\|_{\infty}$, $\|u-\tilde{u}\|_{2}$, the number of iterations and CPU time needed for convergence. We also demonstrate the effectiveness of interpolating the Bellman functions, i.e. $B_{k}(x)$ from Equation (5.1), in our new algorithm on Example 6.1.2.
6.1. Smooth, strictly convex examples
6.1.1. Standard example
The first example that we will look at has the exact solution
$u(x,y)=e^{\frac{1}{2}(x^{2}+y^{2})}$
of the Monge–Ampère equation with a right-hand side given by
$\operatorname{MA}(u)=f(x,y)=(1+x^{2}+y^{2})e^{x^{2}+y^{2}}.$
This example has been used in several previous papers, see for example [5, 12, 18]. Clearly, $u$ is $C^{\infty}$-smooth and strictly convex, so we expect any reasonable solver to give good results for this example.
In fact, the Bellman method produces a sequence of strictly convex approximants, and the interpolation step is never invoked. Comparing the Bellman method to the other methods, all approaches converge to the same solution regardless of mesh size, see Figure 1(a). The Bellman method needs 6–7 iterations to converge, whereas the M2 method requires 40–50 iterations, independent of mesh size. Even if each iteration in our method is computationally a little more demanding than in the M2 method, the overall running time is 5–6 times faster for the Bellman method, see Figure 1(b). Both methods have similar running time complexity ($\sim N^{2.8}$ and $\sim N^{2.9}$, respectively).
[…]
(a) Error in supremum (top) and $\ell^{2}$ norm (bottom). The results for all three algorithms are almost identical and all graphs overlap. The results suggest that $\|u-u_{N}\|\sim N^{-2}$ for the sup-norm as well as the $\ell^{2}$-norm.
(b) Running time (s). The results suggest that the running time is $\sim N^{2.8}$ for the Bellman method, $\sim N^{2.9}$ for M2 method and $\sim N^{4.0}$ for M1.
Figure 1. Standard example: $u(x,y)=e^{0.5(x^{2}+y^{2})}$.
6.1.2. Degenerate example with regularisation
(a) Error in supremum (top) and $\ell^{2}$ norm (bottom). The results for the three algorithms are almost identical and suggest that $\|u-u_{N}\|\sim N^{-2}$, for both norms. The Bellman algorithm without interpolation (Bellman2 in the graph) does not converge.
(b) Running time (s). The results suggest that the running time is $\sim N^{2.7}$ for the Bellman method, $\sim N^{2.8}$ for the M2 method and $\sim N^{4.0}$ for M1.
Figure 2. Degenerate example with regularisation: $u(x,y)=0.5(x-0.5)^{4}+0.1x^{2}+y^{2}$.
(a) Iteration 1
(b) Iteration 2
(c) Iteration 3
(d) Iteration 4
Figure 3. Degenerate example with regularisation: $u(x,y)=0.5(x-0.5)^{4}+0.1x^{2}+y^{2}$.
The Bellman algorithm converges in 9 iterations (mesh size $N=255$). The dark areas in the picture show the region where the computed Hessian $D^{2}u_{k}$ fails to be positive definite. After four iterations, the approximant is everywhere strictly convex and the interpolation step is no longer required.
For this example, the exact solution to the Monge–Ampère equation is
(6.1) $u(x,y)=0.5(x-0.5)^{4}+\epsilon x^{2}+y^{2},$
where ϵ+italic-$\epsilon\in\mathbb{R}_{+}$, and the right-hand side of the equation is
$\ \operatorname{MA}(u)=f(x,y)=12(x-0.5)^{2}+4\epsilon.$
We note that the solution is smooth and strictly convex for each $\epsilon>0$, but the smaller the choice of $\epsilon$, the less convex the exact solution is along the line $x=0.5$. We solve the Monge–Ampère equation for $\epsilon=0.1$ and observe that all three algorithms solve the Monge–Ampère equation successfully.
However, the Bellman algorithm requires the additional interpolation step (step 3) in order to converge, see Figure 2(a). In fact, the solution to the corresponding Poisson equation, $\Delta u=2\sqrt{f}$ fails to be convex when $0\leq\epsilon\lesssim 0.2$, thus the first approximant produced by the Bellman method (as well as M2) is not convex. After a few iterations in the Bellman method, convexity of $u_{k}$ is achieved and the interpolation step is no longer needed, see Figure 3.
[…]
6.2. Mildly singular examples
6.2.1. Trigonometric
On the domain $\Omega=[0,1]^{2}$ we consider the example where the exact solution is given by
$u(x,y)=-\cos{(0.5\pi x)}-\cos{(0.5\pi y)}$
with the corresponding right hand side of the Monge–Ampère equation
$\operatorname{MA}(u)=f(x,y)=(0.5\pi)^{4}\cos{(0.5\pi x)}\cos{(0.5\pi y)}.$
[…]
(a) Error in supremum (top) and $\ell^{2}$ norm (bottom). The results for all three algorithms are almost identical and suggest that $\|u-u_{N}\|\sim N^{-2}$ for both norms. All three graphs overlap in the picture.
(b) Running time (s). The results suggest that the running time is $\sim N^{2.7}$ for the Bellman method and $\sim N^{2.8}$ for the M2, while the running time for M1 is $\sim N^{4.0}$.
Figure 4. Trigonometric: $u(x,y)=-\cos{(0.5\pi x)}-\cos{(0.5\pi y)}$.
6.2.2. Degenerate example
Let us reconsider Example 6.1.2, i.e. Equation (6.1), and remove the regularisation term by setting $\epsilon=0$. In other words, we consider the exact solution
$u(x,y)=0.5(x-0.5)^{4}+y^{2},$
where
$\operatorname{MA}(u)=f(x,y)=12(x-0.5)^{2}.$
Note that the right-hand side vanishes along the line $x=0.5$, and while we can’t expect the Bellman method to produce approximants with positive definite Hessians along this line, nevertheless, all three algorithms successfully tackle the degeneracy and converge to the correct solution, again producing the same results.
As in the strictly convex examples, the Bellman method requires around 10 iterations to converge, regardless of mesh size, but for this example the M2 method struggles, and the number of iterations grows, at a roughly linear rate, as the mesh size $N$ increases, see Table 1. This phenomenon heavily penalises the running time, which in the strictly convex settings was observed to be roughly proportional to $N^{\alpha}$ with $\alpha\approx 2.7$ for both the Bellman and M2 methods.
For this example, the running time of the Bellman method is still around $N^{2.7}$, but for the M2 method, the running time increases to $\sim N^{3.5}$ giving the Bellman method a huge advantage for large mesh sizes. For a $511\times 511$ mesh, the Bellman method ran in just over four minutes, and the method M2 needed almost eleven hours (on a reasonably fast personal computer). Here, the performance of the M1 method was closer to the M2 method, but still about three times slower.
Table 1. Running time for the degenerate example $u=0.5(x-0.5)^{4}+y^{2}$ on an $N\times N$ mesh. We only ran M1 on grid sizes up to $255\times 255$.
Bellman M1 M2
$N$ iter time (s) iter time (s) iter time (s)
31 9 0.1 1990 3.0 260 2.0
63 9 0.7 7947 54.4 452 23.7
127 11 5.7 30 5593055930\,55930 559 867.7 758 269.6
255 10 33.8 115 045115045115\,045115 045 13 182.113182.113\,182.113 182.1 1217 3327.4
511 10 243.1 1859 38 946.938946.938\,946.938 946.9
(a) Error in supremum (top) and $\ell^{2}$ norm (bottom). The results for all three algorithms are once again almost identical and suggest that $\|u-u_{N}\|\sim N^{-2}$ for both norms.
(b) Running time (s). The results suggest that the running time is $\sim N^{2.7}$ for the Bellman method, $\sim N^{3.5}$ for M2 method and $\sim N^{4.0}$ for M1.
Figure 5. Degenerate example: $u(x,y)=0.5(x-0.5)^{4}+y^{2}$.
6.2.3. Constant Monge–Ampère
Let us consider the Dirichlet problem
(6.2) $\begin{cases}\operatorname{MA}(u)\equiv 1&\text{on $\Omega=[-1,1]^{2}$}\\ u\equiv 1&\text{on $\partial\Omega$}\end{cases}$
The exact solution of (6.2) is unknown, and even though the setup looks innocent enough at first glance, this example has shown to be a challenge to most if not all numerical algorithms that have been proposed for the Monge–Ampère equation. The problem is that the boundary conditions force $u$ to be almost constant on lines parallel to and close to the boundary of the square, and the condition $\det(D^{2}u)=1$ then forces the second order derivative in the normal direction to blow up. Consequently, the eigenvalues of the Hessian are neither bounded from above nor away from zero on $\bar{\Omega}$. It follows from results by Błocki [7] that the solution is $C^{1,1}(\Omega)\cap C(\bar{\Omega})$, but by the argument above, we can’t expect higher order regularity of $u$ up to $\partial\Omega$.
Figure 6. Constant MA example. The figure show cross-sections along the domain’s diagonal of the solutions given by the Bellman and M1 algorithm, $u_{B}$ (top) and $u_{M1}$ (bottom), respectively. Both solutions fail to be convex near the corners of the square
Table 2. Minimum value for the computed solution to the constant Monge–Ampère example. Method M1 and M2 produce the same values (to four points of accuracy). The computations with the M2 method were terminated after 10 0001000010\,00010 000 iterations, and the computations with M1 was terminated after 300 000300000300\,000300 000 iterations, marked by * in the table. The results for the wide stencil finite difference discretizations (4th–6th column) are taken from Oberman [26].
$N$ Bellman M1, M2 9pt stencil 17pt stencil 33pt stencil
21 0.2917 0.2892 0.3115 0.2815 0.2839
41 0.2772 0.2734 0.3090 0.2807 0.2732
61 0.2721 0.2682 0.3082 0.2803 0.2711
81 0.2712 0.2655 0.3078 0.2802 0.2704
101 0.2694 0.2639 0.3076 0.2800 0.2700
127 0.2677 0.2626
255 0.2651 *
511 0.2628 *
Table 3. Running time for the constant Monge–Ampère example. Computations marked by * were terminated after 10 0001000010\,00010 000 (for method M2) and 300 000300000300\,000300 000 (for M1) iterations.
Bellman M1 M2
$N$ iter time (s) iter time (s) iter time (s)
$31$ $17$ $0.1$ $2902$ $4.4$ $438$ $6.7$
$41$ $19$ $0.5$ $5086$ $14.3$ $1539$ $23.2$
$61$ $19$ $1.6$ 11 1461114611\,14611 146 $69.9$ $3224$ $150.3$
$81$ $21$ $3.8$ 19 3791937919\,37919 379 $222.5$ $5443$ $546.8$
$101$ $19$ $5.9$ 29 6982969829\,69829 698 $518.3$ $8158$ $1513.7$
$127$ $20$ $10.8$ 46 1524615246\,15246 152 $1297.7$
$255$ $29$ $102.7$ 174 401174401174\,401174 401 20 896.020896.020\,896.020 896.0 * *
$511$ $25$ $870.7$
Running the Bellman algorithm on Equation (6.2) produces a function that (barely) fails to be convex near the corners of $\Omega$, but the same is true for Method M2, see Figure 6. This example was also studied by Dean and Glowinski [10] (using a variational formulation) and Oberman [26] (wide stencil finite difference discretization), all with similar results. In [5], they suggest comparing the different methods using the minimum (i.e. the value at the origin) of the solution. The wide stencils methods from Oberman [26] are known to converge monotonically to a viscocity supersolution of Equation (2.1), and are thus too large. If the solution produced by the Bellman method had been convex, Proposition 3.3 would have implied that it too is too large, but since convexity fails near the corners of the square, we have no guarantee for this. Our method produces a solution which is slightly larger than the M2 and M1 methods, see Table 2.
The two methods used in [5] were observed to have running times of $\sim N^{4.0}$ (M1) and $\sim N^{4.6}$ (M2), respectively (results were reported for grid sizes $21\leq N\leq 141$). When we ran those two methods on grid sizes up to $255\times 255$ (for M1), we observed similar running times.
The M2 method converges very slowly even for moderately large grid sizes, and was terminated after 10,000 iterations when $N\geq 127$. The available results for smaller meshes suggest a running time of $\sim N^{4.6}$. For this example, M1 outperforms M2 and runs at $\sim N^{4.0}$. The Bellman method is an order of magnitude faster still, and clocks in at $\sim N^{3.0}$, see Table 3. Already on a $101\times 101$ grid, the Bellman method runs more than 300 times faster than the M2 method and about 80 times faster than M1.
6.3. Singular examples
6.3.1. Circular degeneracy
(a) Error in supremum and $\ell^{2}$ norm. The errors for Bellman method are worse with $\|u-u_{N}\|_{\infty}\sim N^{-0.8}$ and $\|u-u_{N}\|_{\ell^{2}}\sim N^{-1.0}$. The other methods produce almost identical results with $\|u-u_{N}\|_{\infty}\sim N^{-1.3}$ and $\|u-u_{N}\|_{\ell^{2}}\sim N^{-1.5}$.
(b) Running time (s). The results suggest that the running time is $\sim N^{2.9}$ for the Bellman method, $\sim N^{3.9}$ for the M2 method and ${}\sim N^{4.0}$ for M1.
(c) Difference $u_{B}-u$ (light) and $u_{M2}-u$ (dark), given by Bellman and Method M2, respectively.
(d) Comparing the running time to the achieved accuracy ($\ell^{2}$ norm), the performance of the three methods are fairly similar.
Figure 7. Circular Degeneracy:
$u(x,y)=\frac{1}{2}(\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}-0.2)^{+}$.
We consider
(6.3) $u(x,y)=0.5(\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}-0.2)^{+},$
which would be a solution to the Monge–Ampère equation with the following right-hand side
$\operatorname{MA}(u)=f(x,y)=\frac{(\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}-0.2)^{+}}{\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}}.$
In this example, $u\in C^{1,1}$, but $f=\operatorname{MA}(u)$ vanishes on a disc centred at $(0.5,0.5)$ with radius $r=0.2$. Here, both the M1 and the M2 methods outperform the Bellman algorithm, see Figure 7(a). As expected, the Bellman algorithm struggles on large areas where the exact solution is not strictly convex, that is on the aforementioned disc. This is the area where the solutions $u_{\text{B}}$ and $u_{M2}$ deviate the most from the exact solution $u$, see Figure 7(c). In fact, we see that the error in the supremum and $\ell^{2}$ norms are proportional to $N^{-0.8}$ and $N^{-1.0}$, respectively, for the Bellman algorithm, whereas they are around $N^{-1.3}$ and $N^{-1.5}$ for the other methods.
While the accuracy of the Bellman method is notably worse than the other methods, it seems like all methods still converge to the exact solution as the mesh size $N$ tends to $\infty$, but the rate of convergence, for all methods, is slower than for the strictly convex examples, see Figure 7(a).
Even though the Bellman method gives worse results, it still outperforms the other methods with respect to running time. The measured time complexity for the Bellman method is similar to the strictly convex examples ($\sim N^{2.9}$), whereas the running time of the M2 method is $\sim N^{3.9}$ , resulting in a running time that is about 70 times slower than the Bellman method for a $511\times 511$ mesh. The number of iterations required is 6–14 for the Bellman method, where Method M2 needs an increasing number of iterations as the number of mesh points increases (1092 for the $511\times 511$ mesh). The M1 method runs even slower. If we look at accuracy as a function of computational time (see Figure 7(d)), we see that all three methods perform fairly similarly.
Remark.
We would like to remark that the Bellman algorithm fails to converge when the right-hand side vanishes on the domain, i.e. $f\equiv 0$ on $\Omega$. One example of such a function could be $u(x,y)=|x|$. Here, the solution given by our method is the solution to the Poisson equation that is the first iteration. Method M2 converges. However, the convergence is very slow, and the running time grows from 1 min to 6 hours as the mesh size increases, and with diminished accuracy, as the method is terminated after 10 000 iterations.
6.3.2. Unbounded Monge–Ampère
We investigate the convergence for an unbounded right-hand side. For $\Omega=[0,1]^{2}$ and
$u(x,y)=-\sqrt{2-x^{2}-y^{2}}.$
The right-hand side of the equation is
$\operatorname{MA}(u)=f(x,y)=\frac{2}{(2-x^{2}-y^{2})^{2}}$
and we note that $f$ is unbounded near the point $(1,1)$. All three algorithms struggle, and only achieve an accuracy of order $N^{-0.5}$ (measured in sup-norm), which is the same as observed by Benamou et al. [5] for M1 and M2. Most of the error is concentrated near the problematic point $(1,1)$, and if convergence is measured by average $\ell^{2}$ norm instead, the accuracy is $N^{-1.5}$. All algorithms run slowly, and the Bellman method and M2 runs at roughly the same speed, while M1 is about 20 times slower.
Curiously enough, the Bellman method seems to enjoy a relative speed-up as well as an improvement in accuracy for larger meshes, see Figure 8.
If we instead do the computations on a slightly smaller square, e.g. $[0,0.99]^{2}$, the three methods run at full accuracy ($\sim N^{-2}$), and the Bellman method runs 2–3 times faster than M2.
(a) Error in supremum (top) and $\ell^{2}$ norm (bottom). The results for all three algorithms are fairly close, but not identical, and suggest that $\|u-u_{N}\|_{\infty}\sim N^{-0.5}$ and $\|u-u_{N}\|_{\ell^{2}}\sim N^{-1.5}$.
(b) Running time (s). The results suggest that the running time is $\sim N^{3.0}$ for the Bellman method, $\sim N^{3.1}$ for M2 and $\sim N^{4.0}$ for M1.
(c) Comparing the running time to achieved accuracy ($\ell^{2}$ norm), the performance of the three methods are fairly similar, with the Bellman method winning for very fine meshes.
Figure 8. Unbounded MA: $u(x,y)=-\sqrt{2-x^{2}-y^{2}}$.
7. Conclusions
[…]
Similar performance was also observed in cases where the right hand side is mildly degenerate, such as the set where $f=0$ is at most one dimensional. In this setting, our method requires a few more iterations to converge (typically 10–15), but this only has a minor effect on the running time. In comparison, the M2 method converges a lot slower for this class of examples, typically needing more and more iterations as the number of grid points increases, resulting in running times that vary between $N^{2.8}$ and $N^{3.6}$.
The net effect is that our method runs 3–10 times faster than the M2 method on smooth strictly convex examples, but as much as 20–100 (or more) times faster for the mildly degenerate cases. The M1 method is even slower, and has a running time proportional to around $N^{4.0}$ for most examples. With the exception of the constant Monge–Ampère example, M1 runs substantially slower than M2.
In the most challenging examples, when $f$ vanishes on an open set or when $f$ is unbounded, the Bellman method begins to struggle. While the running time and the number of iterations is similar to the less challenging examples, the accuracy becomes worse resulting in supremum norm error between $N^{-0.8}$ and $N^{-0.5}$. In these examples, the errors are mostly concentrated near the zero set (or the unbounded locus) of $f$, and measuring the error in $\ell^{2}$ norm gives slightly better results, with an error between $N^{-1.0}$ and $N^{-1.5}$.
On the other hand, other published methods struggle with these more challenging examples as well. Method M2 outperforms the Bellman method in the circular degeneracy example (Example 6.3.1), at the cost of very slow running times. In all the other examples that we have considered, the accuracy of the Bellman method, M2 method and M1 are very similar, indicating that all methods converge to the optimal discretization of the underlying PDE on the given mesh, and in almost all examples the Bellman method runs considerably faster than the remaining methods.

### v3

@@
 $\mu_{u}(E)=\operatorname{Leb}(\partial u(E))$
-for any Borel subset $E\subset\Omega$, where $\operatorname{Leb}$ denotes the Lebesgue measure, $\partial u(E)=\cup_{x\in E}\partial u(x)$, and $\partial u(x)$ is the set of all slopes of the supporting hyperplanes to the graph of $u$ at $x$.
+for any Borel subset $E\subset\Omega$, where $\operatorname{Leb}$ denotes the Lebesgue measure, $\partial u(E)=\cup_{x\in E}\partial u(x)$, and $\partial u(x)$ is the subgradient of $u$, i.e.
+the set of all slopes of the supporting hyperplanes to the graph of $u$ at $x$.
 If $u\in C^{2}$ and $D^{2}u$ is positive semi-definite, then $\mu_{u}(E)=\det(D^{2}u)$.
@@
 A complete overview is outside the scope of this paper, but for example, it follows from results by Aleksandrov [1] and Bakelman [3] from the 1950’s that if $\Omega$ is assumed to be a bounded strictly convex domain, $f$ is a finite Borel measure on $\Omega$, and $\phi\in C(\partial\Omega)$, then (2.1) has a unique convex solution $u$ (in the sense of Aleksandrov) with $u\in C(\bar{\Omega})$.
-Furthermore, if $\Omega$ is strictly convex (or uniformly convex in the unbounded setting) with $C^{k+2,\alpha}$ boundary, $0<\phi\in C^{k+2,\alpha}(\partial\Omega)$ and $f\in C^{k,\alpha}(\bar{\Omega})$ for some integer $k\geq 2$ and some $10<\alpha<$ 1, then the Dirichlet problem (2.1) has a unique convex solution $u\in C^{k+2,\alpha}(\bar{\Omega})$, see [9, 21, 22].
+Furthermore, if $\Omega$ is strictly convex (or uniformly convex in the unbounded setting) with $C^{k+2,\alpha}$ boundary, $0<\phi\in C^{k+2,\alpha}(\partial\Omega)$ and $f\in C^{k,\alpha}(\bar{\Omega})$ for some integer $k\geq 2$ and some $0<\alpha<1$, then the Dirichlet problem (2.1) has a unique convex solution $u\in C^{k+2,\alpha}(\bar{\Omega})$, see [9, 21, 22].
 If we merely assume that $\Omega$ is convex, not necessarily strictly convex, the situation is a little less straight-forward, and we need to assume that $\phi$ is convex on all line segments in $\partial\Omega$ to guarantee existence of solutions to (2.1).
@@
 As far as we are aware, the first approach was due to Oliker and Prussner (1988) [28] who proposed a direct discretization of the Aleksandrov interpretation of the Monge–Ampère operator.
-Since then, a number of other approaches to solving Monge–Ampère equations have been tried: finite difference discretization, often using wide stencils to obtain stable and consistent methods [5, 6, 11, 12, 26], variational methods [4, 19], fixpoint methods [5, 18] as well as finite elements [20, 23] only to mention a few.
+Since then, a number of other approaches to solving Monge–Ampère equations have been tried: finite difference discretization, often using wide stencils to obtain stable and consistent methods [5, 6, 11, 12, 27], variational methods [4, 19], fixed-point methods [5, 18] as well as finite elements [20, 23] only to mention a few.
+After the preparation of an earlier draft of this paper, we learned that a similar method as ours (for periodic boundary values) has been previously proposed by Loeper and Rapetti [26].
+Their approach is somewhat different than ours, and derive their method as a Newton type iteration for the non-linear differential operator.
 3. Bellman’s principle
 Our method relies on the following linear algebra result, sometimes called Bellman’s principle.
-In the setting of complex Monge–Ampère equations, this idea was introduced by Gaveau [15] and has later found several theoretical applications.
-We will formulate the result in the real setting and for convenience, we provide a proof of Bellman’s principle, similar to the approach taken by Wiklund [29].
+In the setting of complex Monge–Ampère equations, this idea was introduced by Gaveau [15] and has later found several theoretical applications, both in the real and the complex setting.
+We will formulate the result in the real setting and for convenience of the reader, we provide a proof of Bellman’s principle.
 Theorem 3.1.
@@
 $(\det M)^{1/n}\leq\frac{1}{n}\inf_{B\in\mathcal{B}}\operatorname{tr}(BM).$
-To this end we observe that for any symmetric and positive definite matrix $B$, we have
+To this end, we observe that for any symmetric and positive definite matrix $B$, we have
 $\begin{split}\det(BM-\lambda I)&=\det(B^{1/2}B^{1/2}MB^{1/2}B^{-1/2}-\lambda I)\\ &=\det(B^{1/2})\det(B^{1/2}MB^{1/2}-\lambda I)\det(B^{-1/2})\\ &=\det(B^{1/2}MB^{1/2}-\lambda I).\end{split}$
@@
 (3.2) $\begin{cases}\operatorname{tr}(B(x)\,D^{2}u(x))=n\,\big(f(x)\big)^{1/n},&x\in\Omega,\\ u(x)=\phi(x),&x\in\partial\Omega,\end{cases}$
-for which there are highly efficient numerical algorithms. Let us mention that with this approach we have the following monotonicity result.
-Proposition 3.3.
-Let $\Omega\subset\mathbb{R}^{n}$ be a convex domain with $C^{2,\alpha}$ boundary, $\phi\in C^{2,\alpha}(\partial\Omega)$, $f^{1/n}\in C^{0,\alpha}(\Omega)$ and let $B\in C^{0,\alpha}(\Omega)$ be a $\mathcal{B}$-valued function on $\Omega$.
-Assume that the solution to the above Dirichlet problem (Eq.
-(3.2)) is convex.
-Then $u\geq u_{\text{MA}}$ where $u_{\text{MA}}$ denotes the solution to the corresponding Monge–Ampère equation (2.1).
-Proof.
-By standard regularity theory for elliptic PDEs, $u\in C^{2,\alpha}$, so by Corollary 3.2, $\operatorname{MA}(u)\leq f$.
-Hence, by the comparison principle for the real Monge–Ampère operator (see for example [24, Theorem 3.21]), it follows that $u\geq u_{\text{MA}}$.
-∎
+for which there are highly efficient numerical algorithms.
 Unfortunately, it seems like there are no known effective characterizations of when the solution to Equation (3.2) is convex, even in the case when $B\equiv I$, i.e.
 for Poisson’s equation.
-Our numerical experiments suggest that our proposed algorithm (as described in Section 5) produces a decreasing sequence of approximants converging to the solution of (2.1), at least once one of the approximants is convex.
-So far, we have not been able to prove this in full generality, but under suitable technical conditions and the assumption that the method produces a sequence of strictly convex approximants, we can show convergence, see Theorem 4.1.
+Under suitable technical conditions and the assumption that our method produces a sequence of strictly convex approximants, we can show convergence, see Theorem 4.1.
 We should point out that there is no proof of convergence for most methods that have been proposed for the Monge–Ampère equation.
@@
 Notable methods with a proof of convergence are the direct discretization of the Aleksandrov construction proposed in [28] and the wide stencil finite difference discretizations [27] which are also known to converge (to a viscosity solution).
-In the remainder of the paper, we will restrict our attention to the two-dimensional case, $n=2$, but we expect the method to be competitive also in higher dimensions.
 4. Proof of convergence
@@
 Theorem 4.1.
-Let $\alpha\in(0,1)$ and let $\Omega\subset\mathbb{R}^{2}$ be a strictly convex, bounded, $C^{2,\alpha}$ domain, $\phi\in C^{2,\alpha}(\bar{\Omega})$ and $f\in C^{0,\alpha}(\bar{\Omega})$ where $f>0$ on $\bar{\Omega}$.
-Assume that there exist strongly convex solutions $\{u_{k}\}_{k\geq 0}\subset C^{2,\alpha}(\bar{\Omega})$ to the following linear Dirichlet problems
-(4.1) $\begin{cases}\operatorname{tr}(B_{k}(x)D^{2}u_{k}(x))=2\sqrt{f(x)},&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega,\end{cases}$
-where $B_{k}=\sqrt{\det D^{2}u_{k}}\,(D^{2}u_{k})^{-1}$ for $k\geq 1$ and $B_{0}$ is a $\mathcal{B}$-valued function with coefficients $(B_{0})_{ij}\in C^{0,\alpha}(\bar{\Omega})$.
-Moreover, denote by $\lambda_{k},\Lambda_{k}$ the eigenvalues of $D^{2}u_{k}$ and assume that there exists $\epsilon>0$ such that $\epsilon\leq\lambda_{k},\Lambda_{k}$ for all $k$.
-Then, there exists a subsequence of the solutions $\{u_{k}\}_{k\geq 0}$ converging uniformly to a convex function $\tilde{u}\in C^{1,1}(\bar{\Omega})$ solving the Monge–Ampère equation (2.1).
+Let $\alpha\in(0,1)$ and let $\Omega\subset\mathbb{R}^{n}$ be a strictly convex, bounded, $C^{2,\alpha}$ domain, $\phi\in C^{2,\alpha}(\bar{\Omega})$ and $f\in C^{0,\alpha}(\bar{\Omega})$ with strictly positive values on $\bar{\Omega}$.
+Assume that there exist solutions $\{u_{k}\}_{k\geq 0}\subset C^{2,\alpha}(\bar{\Omega})$ to the following linear Dirichlet problems
+(4.1) $\begin{cases}\operatorname{tr}(B_{k}(x)D^{2}u_{k}(x))=nf^{1/n}(x),&x\in\Omega\\ u_{k}(x)=\phi(x),&x\in\partial\Omega\\ u_{k}(x)\text{ is strongly convex}.\end{cases}$
+where $B_{k}=(\det{D^{2}u_{k-1}})^{1/n}(D^{2}u_{k-1})^{-1}=\operatorname{cof}D^{2}u_{k-1}$ for $k\geq 1$ and $B_{0}$ be a $\mathcal{B}$-valued function with coefficients $(B_{0})_{ij}\in C^{0,\alpha}(\bar{\Omega})$.
+Moreover, denote by $\lambda_{\text{min}}^{k}$ the smallest eigenvalue of $D^{2}u_{k}$ and assume that $\{\lambda_{\min}^{k}\}_{k\geq 0}$ is uniformly bounded away from zero, i.e.
+there exists $\epsilon>0$ such that $\lambda_{\text{min}}^{k}\geq\epsilon$ for every $k\geq 0$.
+Then, the sequence of solutions $\{u_{k}\}_{k\geq 0}$ converges uniformly to a convex function $\tilde{u}\in C^{1,1}(\bar{\Omega})$ solving the Monge–Ampère equation (2.1).
 Before proving the above theorem, let us first discuss the requirement of the existence of solutions to (4.1) and their strong convexity.
 Under the regularity assumptions stated above on the domain $\Omega$, the functions $f$ and $\phi$, and the coefficients $(B_{0})_{ij}$, standard theory of linear elliptic PDEs (see Gilbarg and Trudinger [16, Theorem 6.14]) provides existence and uniqueness of a solution $u_{0}\in C^{2,\alpha}(\bar{\Omega})$.
-Now, let us investigate solvability and well-definiteness of the next Dirichlet problem in the sequence.
+Next, let us investigate solvability and well-definiteness of the next Dirichlet problem in the sequence.
 We would like to generalise this argument for any $k\geq 0$.
 To this end, assume the existence of the solution $u_{k}\in C^{2,\alpha}(\bar{\Omega})$.
-Observe that $B_{k+1}\coloneqq\sqrt{\det D^{2}u_{k}}(D^{2}u_{k})^{-1}$ is not necessarily well-defined, nor is the next differential operator $L_{k+1}(v)\coloneqq\operatorname{tr}(B_{k+1}D^{2}v)$ elliptic.
+Observe that $B_{k+1}\coloneqq(\det D^{2}u_{k})^{1/n}\,(D^{2}u_{k})^{-1}$ is not necessarily well-defined, nor is the next differential operator $L_{k+1}(v)\coloneqq\operatorname{tr}(B_{k+1}D^{2}v)$ strictly elliptic.
 Consequently, the solution to the corresponding Dirichlet problem is not guaranteed.
-However, strong convexity of $u_{k}$ renders the next differential operator $L_{k+1}$ well-defined and elliptic.
-To this end, note that $D^{2}u_{k}\in C^{0,\alpha}(\bar{\Omega})$ and thus $\det(D^{2}u_{k})\in C^{0,\alpha}(\bar{\Omega})$.
-Hence, both eigenvalues are bounded from above.
-Now, if $\epsilon\leq\lambda_{k},\Lambda_{k}$ it follows that $D^{2}u_{k}$ is positive definite and hence invertible with a positive determinant.
-Thus $B_{k+1}$ is well-defined and by construction it is also positive definite.
+However, strong convexity of $u_{k}$ renders the next differential operator $L_{k+1}$ well-defined and strictly elliptic.
+Indeed, first note that from $\epsilon\leq\lambda_{\text{min}}^{k}$ it follows that $D^{2}u_{k}$ is positive definite and hence invertible.
+Recall that $B_{k+1}\coloneqq(\det D^{2}u_{k})^{1/n}(D^{2}u_{k})^{-1}$, by the above $B_{k+1}$ is well-defined and positive definite with $\det B_{k+1}=1$.
+We can use the Corollary to Bellman’s principle (Corollary 3.2) to conclude that
+$n\epsilon\leq n(\det D^{2}u_{k})^{1/n}=\operatorname{tr}{B_{k+1}D^{2}u_{k}}\leq\operatorname{tr}{B_{k}D^{2}u_{k}}=nf^{1/n}\text{ on }\bar{\Omega}.$
+Thus,
+(4.2) $\epsilon^{n}\leq\det D^{2}u_{k}\leq f\text{ on }\bar{\Omega}.$
+Moreover, existence of a uniform lower bound of the eigenvalues $\lambda^{k}_{min}$ implies existence of a uniform upper bound for the largest eigenvalues denoted by $\lambda^{k}_{max}$.
+Indeed, Eq.
+(4.2) yields
+$\epsilon\leq\lambda^{k}_{min}\leq\lambda^{k}_{max}\leq f/\epsilon^{n-1},\qquad\text{for $k\geq 0$.}$
+By diagonalising $B_{k+1}$ we have that the smallest and the largest eigenvalues are uniformly bounded as follows
+$\begin{split}\gamma_{min}^{k}&=(\det D^{2}u_{k})^{1/n}/\lambda_{max}^{k}\geq(\epsilon^{n})^{1/n}/(f/\epsilon^{n-1})\geq\epsilon/\|f\|_{C(\bar{\Omega})}\text{ and }\\ \gamma_{max}^{k}&=(\det D^{2}u_{k})^{1/n}/\lambda_{min}^{k}\leq\|f^{1/n}\|_{C(\bar{\Omega})}/\epsilon.\end{split}$
+From the lower bound of $\gamma_{min}^{k}$ follows strict ellipticity of $L_{k+1}$.
 Furthermore, the regularity of the solution, i.e.
-$u_{k}\in C^{2,\alpha}(\bar{\Omega})$ results in $(B_{k})_{ij}\in C^{0,\alpha}(\bar{\Omega})$.
-Thus $L_{k+1}$ is elliptic with the necessary regularity of coefficients for the corresponding Dirichlet problem to have a unique solution $u_{k+1}\in C^{2,\alpha}(\bar{\Omega})$ (by another application of Theorem 6.14 in Gilbarg and Trudinger).
+$u_{k}\in C^{2,\alpha}(\bar{\Omega})$ results in $(B_{k+1})_{ij}\in C^{0,\alpha}(\bar{\Omega})$ as $(B_{k+1})_{ij}$ are obtained by adding, subtracting and multiplying terms $\partial_{ij}^{2}u_{k}$.
+Thus $L_{k+1}$ is strictly elliptic with the necessary regularity of coefficients for the corresponding Dirichlet problem to have a unique solution $u_{k+1}\in C^{2,\alpha}(\bar{\Omega})$ (by another application of Theorem 6.14 in Gilbarg and Trudinger).
 Proof.
-We begin with defining $\nu_{k}\coloneqq\det D^{2}u_{k}(x)$ and noting that $\{u_{k}\}_{k\geq 0}$ are solutions to the following non-linear Dirichlet problems:
+We begin with showing monotonicity and hence convergence of the sequence of solutions $\{u_{k}\}_{k\geq 0}$.
+Let us define $\nu_{k}\coloneqq\det D^{2}u_{k}(x)$ and $\eta_{k}\coloneqq u_{k+1}-u_{k}$. Then
+$\operatorname{tr}B_{k+1}D^{2}\eta_{k}$ $=\operatorname{tr}B_{k+1}D^{2}u_{k+1}-\operatorname{tr}B_{k+1}D^{2}u_{k}$
+$=nf^{1/n}-\operatorname{tr}(\det D^{2}u_{k})^{1/n}\,(D^{2}u_{k})^{-1}D^{2}u_{k}$
+$=nf^{1/n}-n(\det D^{2}u_{k})^{1/n}=nf^{1/n}-n(\nu_{k})^{1/n}.$
+Hence $\eta_{k}$ solves the following linear elliptic PDE, the coefficients of which are elements of the matrix $B_{k+1}\coloneqq\operatorname{cof}D^{2}u_{k}$,
+$\begin{cases}\operatorname{tr}B_{k+1}D^{2}\eta_{k}=n(f)^{1/n}-n(\nu_{k})^{1/n}\text{ in }\Omega,\\ \eta_{k}=0\text{ on }\partial\Omega.\end{cases}$
+Since the operator $L_{k+1}(v)\coloneqq\operatorname{tr}B_{k+1}(D^{2}v)$ is strictly elliptic and Eq.
+(4.2)) gives $n(f)^{1/n}-n(\nu_{k})^{1/n}\geq 0$, the maximum principle for linear elliptic PDEs ([16, Theorem 3.1]) yields
+$u_{k}-u_{k-1}=\eta_{k}\leq\max_{\bar{\Omega}}\eta_{k}=\max_{\partial\Omega}\eta_{k}=0,$
+or in other words, for every $k\geq 0$.
+(4.3) $u_{k}\leq u_{k-1}\text{ on }\bar{\Omega}.$
+It follows from the above and from the comparison principle for the real Monge–Ampère operator (see for example [24, Theorem 3.21]) that $\{u_{k}\}_{k\geq 0}$ is a decreasing sequence bounded from below by $u_{MA}$, the solution to the corresponding Monge–Ampère equation (2.1).
+Hence, the sequence converges to some function $\tilde{u}\coloneqq\lim_{k\to\infty}u_{k}$.
+Moreover, for every $k\geq 0$ and $\psi\in C^{\infty}_{0}(\Omega)$
+$\int_{\Omega}(\partial_{ij}^{2}u_{k+1})\psi dV=\int_{\Omega}u_{k}(\partial_{ij}^{2}\psi)dV.$
+The monotone convergence theorem implies that all second partial derivatives $\{\partial_{ij}^{2}u_{k}\}_{k}$ converge weak-star.
+Hence, we can define a limit function $\nu\coloneqq\mathop{w^{*}\text{-}\lim}_{k\to\infty}\nu_{k}$.
+Since $0<\epsilon^{n}\leq\nu,\nu_{k}\leq f$ for all $k\geq 0$ (Eq.
+(4.2)), we can interpret $\nu_{k}$ and $\nu$ as finite Borel measures.
+We note that $\{u_{k}\}_{k\geq 0}$ are Aleksandrov solutions to the following non-linear Dirichlet problems:
 (4.4) $\begin{cases}\det D^{2}u_{k}(x)=\nu_{k}(x),&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega,\\ u_{k}\text{ strongly convex.}\end{cases}$
-Moreover, the existence of a uniform lower bound of the eigenvalues $\lambda_{k},\Lambda_{k}$ implies the existence of a uniform upper bound.
-Indeed, recall that since $B_{k+1}=\sqrt{\det D^{2}u_{k}}\,(D^{2}u_{k})^{-1}$, it is positive definite with $\det B_{k+1}=1$, for every $k\geq 0$, and we can use the Corollary to Bellman’s principle (Corollary 3.2) to conclude that
-(4.3) $2\epsilon\leq 2\sqrt{\det D^{2}u_{k}}=\operatorname{tr}{B_{k+1}D^{2}u_{k}}\leq\operatorname{tr}{B_{k}D^{2}u_{k}}=2\sqrt{f}\text{ on }\bar{\Omega}.$
-Thus,
-(4.4) $\epsilon\leq\lambda_{k},\Lambda_{k}\leq\|\sqrt{f}\|_{C(\bar{\Omega})}/\epsilon,\qquad\text{for $k\geq 0$.}$
-Consequently, $\{\nu_{k}\}_{k}$ can be interpreted as finite Borel measures with $\nu_{k}(\Omega)\leq(\|f\|_{C(\Omega)}/\epsilon)^{2}$ and $\{u_{k}\}_{k}$ as Aleksandrov solutions to (4.2).
-Moreover, there exists a finite Borel measure $\nu$ and a subsequence $\{\nu_{k}\}_{k\geq 0}$ (after re-enumeration) such that $\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}\nu_{k}$.
-By compactness of Aleksandrov solutions to the Dirichlet problem (4.2) ([24, Theorem 3.35]), we can conclude that there exists $\tilde{u}\in C(\bar{\Omega})$ and a subsequence $\{u_{k}\}_{k\geq 0}$ (after re-enumeration) such that $u_{k}\to\tilde{u}$ uniformly and $\tilde{u}$ satisfies the following
+Next, we use compactness of Aleksandrov solutions to the Dirichlet problem (4.4) ([24, Theorem 3.35]) to conclude that the limit function $\tilde{u}\coloneqq\lim_{k\to\infty}u_{k}\in C(\bar{\Omega})$ and satisfies the following
 $\begin{cases}\det D^{2}\tilde{u}(x)=\nu(x),&x\in\Omega,\\ \tilde{u}(x)=\phi(x),&x\in\partial\Omega,\\ \tilde{u}\text{ convex.}\end{cases}$
+Since the limit function is continuous and the sequence monotone, Dini’s theorem gives that the convergence is uniform.
 It remains to show that $\tilde{u}$ solves the initial Monge–Ampère equation (2.1) by showing that $f=\nu$ as measures.
-To this end, we observe that all partial derivatives $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ are uniformly bounded from above and below.
-Indeed, we can use Rayleigh quotients to conclude that $\{(D^{2}u_{k})_{ii}\}_{k\geq 0}$ are bounded by the eigenvalues.
-The existence of uniform lower and upper bounds follows from Equation (4.4).
-Now, using $\det D^{2}u_{k}=(D^{2}u_{k})_{ii}(D^{2}u_{k})_{jj}-(D^{2}u_{k})^{2}_{ij}$ and Equation (4.3), we see that $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ are also uniformly bounded.
-Thus we can choose a subsequence $\{u_{k}\}_{k\geq 0}$ (after re-enumeration) such that $u_{k}\to\tilde{u}$ uniformly and $\{(D^{2}u_{k})_{ii}\}_{k\geq 0}$, $\{(D^{2}u_{k})_{jj}\}_{k\geq 0}$ and $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ have a weak limit.
-To show that $f=\nu$ in measures, we recall Equation (4.1) and $\operatorname{tr}(B_{k+1}D^{2}u_{k})=2\sqrt{\det(D^{2}u_{k})}\coloneqq 2\sqrt{\nu_{k}}$ and establish the following
-$2\sqrt{f}$ $=\operatorname{tr}(B_{k}D^{2}u_{k})=\operatorname{tr}(B_{k}-B_{k+1})D^{2}u_{k}+\operatorname{tr}(B_{k+1}D^{2}u_{k})$
-$=\underbrace{\operatorname{tr}(B_{k}-B_{k+1})D^{2}u_{k}}_{(*)}+\underbrace{2\sqrt{\nu_{k}}}_{(**)}.$
-Furthermore, recall that $B_{k}=\sqrt{\det D^{2}u_{k-1}}(D^{2}u_{k-1})^{-1}$, so Cramer’s rule gives
-$\begin{split}(B_{k})_{11}&=(D^{2}u_{k-1})_{22}/\sqrt{\det D^{2}u_{k-1}},\\ (B_{k})_{22}&=(D^{2}u_{k-1})_{11}/\sqrt{\det D^{2}u_{k-1}},\\ (B_{k})_{12}&=-(D^{2}u_{k-1})_{21}/\sqrt{\det D^{2}u_{k-1}}.\end{split}$
-Consequently, $(*)\xrightarrow{w^{*}}0$ as $k\to\infty$ since the denominators are uniformly bounded away from 00 by (4.3).
-Furthermore ()(**)( ∗ ∗ ) is also uniformly bounded by Equation (4.3).
-This yields
-$f-\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(f-\nu_{k})=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(\sqrt{f}-\sqrt{\nu_{k}})(\sqrt{f}+\sqrt{\nu_{k}})=0.$
-Since $u_{k}$ converge uniformly and all the second-order derivatives are uniformly bounded, it follows from the Arzelà–Ascoli theorem that $\tilde{u}\in C^{1,1}(\bar{\Omega})$. ∎
+To this end, let us note the following
+(4.5) $\begin{split}n(f)^{1/n}&=\operatorname{tr}(B_{k+1}D^{2}u_{k+1})=\operatorname{tr}B_{k+1}(D^{2}u_{k+1}-D^{2}u_{k})+\operatorname{tr}B_{k+1}D^{2}u_{k}\\ &=\underbrace{\operatorname{tr}B_{k+1}(D^{2}u_{k+1}-D^{2}u_{k})}_{\coloneqq\tau}+n(\underbrace{\nu_{k}}_{\xrightarrow[]{w^{*}}\nu})^{1/n}.\end{split}$
+Hence, it remains to show that $\tau\to 0$ weak-star.
+Since all second partial derivatives $\partial_{ij}^{2}u_{k}$ converge weak-star, it suffices to uniformly bound terms $(B_{k})_{ij}$.
+To this end, we first define bilinear forms $(x,y)_{k}\coloneqq(B_{k}x,y)$ for all $x,y\in\mathbb{R}^{n}$.
+From Rayleigh quotients, we conclude that $\gamma^{k}_{min}\leq\|x\|^{2}_{k}/\|x\|^{2}\leq\gamma^{k}_{max}$ for all $x\in\mathbb{R}^{n}$ and for all $k\geq 0$.
+Let $\{e_{m}\}_{1\leq m\leq n}$ denote the standard basis in $\mathbb{R}^{n}$ and use the polarization identity to obtain
+$(B_{k})_{ij}=(e_{j},e_{i})_{k}=(1/4)(\|e_{j}+e_{i}\|^{2}_{k}-\|e_{j}-e_{i}\|^{2}_{k}).$
+Recalling Equation (4), it follows that
+(4.6) $\epsilon/\|f\|_{C(\bar{\Omega})}\leq\gamma_{min}^{k}\leq(B_{k})_{ij}\leq\gamma_{max}^{k}\leq\|f^{1/n}\|_{C(\bar{\Omega})}/\epsilon,$
+for all $1\leq i,j\leq n$ and $k\geq 0$. Thus, we have that
+$\tau\coloneqq\operatorname{tr}B_{k+1}(D^{2}u_{k+1}-D^{2}u_{k})\xrightarrow[]{w^{*}}0.$
+By the above and Equation (4.5) we have
+$f-\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(f-\nu_{k})=(f^{1/n}-\nu_{k}^{1/n})\sum_{j=0}^{n-1}(-1)^{j}f^{(n-1-j)/n}\nu_{k}^{j/n}=0.$
+Thus, $f=\nu$ as measures. Since all second partial derivatives of $\{u_{k}\}_{k\geq 0}$ are bounded, the Arzelà–Ascoli theorem gives that in fact $\tilde{u}\in C^{1,1}(\bar{\Omega})$. ∎
 5. Description of our algorithm
+In the remainder of the paper, we will restrict our attention to the two-dimensional case, $n=2$, but we expect the method to be competitive also in higher dimensions.
 Theorem 4.1 provides the idea for our numerical method, which we name Bellman’s algorithm and describe below.
@@
 For simplicity of the implementation, we only consider the case where $\Omega$ is a bounded rectangle in $\mathbb{R}^{2}$, despite the fact that these domains are not strictly convex, and the theoretical framework for solving Monge–Ampère equations on rectangles is less well-developed.
-Next, we construct a sequence of functions $u_{k}$ where $u_{k}$ solves the discretized version of the following Dirichlet problem, using a narrow stencil second order finite difference solver:
+Next, we construct a sequence of functions $u_{k}$ where $u_{k}$ solves the discretized version of the following Dirichlet problem, using a central difference second order finite difference solver:
 (5.1) $\begin{cases}\operatorname{tr}(B_{k}(x)D^{2}u_{k}(x))=2\sqrt{f(x)},&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega.\end{cases}$
@@
     We loop over the mesh a second time and replace $B_{k}(x)$ at the marked points with the (normalized to determinant 1) convex combination of the computed Bellman matrices corresponding to the nearest grid points in horizontal and vertical direction for which $D^{2}u_{k-1}$ was positive definite.
-We terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or when the number of iterations exceeds 10 0001000010\,00010 000.
+We terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or when the number of iterations exceeds 10 00010\,000.
 If the solution of the original Monge–Ampère equation is strictly convex, we expect the constructed functions $u_{k}$ to be strictly convex as well, at least when $k$ sufficiently large, but it can happen that convexity fails in the first few steps – see Example 6.1.2 – even when the final solution is strictly convex.
@@
 We should note that this is computationally expensive, especially in our NumPy implementation where explicit loops are slow, so M1, and to some extent our proposed method, are the ones that would benefit the most from an optimized implementation.
+Here, it would be natural to update the whole matrix of functions values at once, using (5.4) instead of doing it element by element, but apparently this destroys the convergence to the actual solution of the Monge–Ampère equation.
 In our comparison we chose the same termination criterion for all methods, i.e.
-to terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or the number of iterations exceeds 10 0001000010\,00010 000 (for M1, we allow up to 300 000300000300\,000300 000 iterations, where one iteration means updating each $u_{ij}$ once).
+to terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or the number of iterations exceeds 10 00010\,000 (for M1, we allow up to 300 000300\,000 iterations, where one iteration means updating each $u_{ij}$ once).
 6. Numerical experiments
@@
 (6.1) $u(x,y)=0.5(x-0.5)^{4}+\epsilon x^{2}+y^{2},$
-where ϵ+italic-$\epsilon\in\mathbb{R}_{+}$, and the right-hand side of the equation is
+where $\epsilon\in\mathbb{R}_{+}$, and the right-hand side of the equation is
 $\operatorname{MA}(u)=f(x,y)=12(x-0.5)^{2}+4\epsilon.$
@@
 Bellman M1 M2
-$N$ iter time (s) iter time (s) iter time (s)
+NN iter time (s) iter time (s) iter time (s)
 31 9 0.1 1990 3.0 260 2.0
 63 9 0.7 7947 54.4 452 23.7
-127 11 5.7 30 5593055930\,55930 559 867.7 758 269.6
-255 10 33.8 115 045115045115\,045115 045 13 182.113182.113\,182.113 182.1 1217 3327.4
-511 10 243.1 1859 38 946.938946.938\,946.938 946.9
+127 11 5.7 30 55930\,559 867.7 758 269.6
+255 10 33.8 115 045115\,045 13 182.113\,182.1 1217 3327.4
+511 10 243.1 1859 38 946.938\,946.9
 (a) Error in supremum (top) and $\ell^{2}$ norm (bottom). The results for all three algorithms are once again almost identical and suggest that $\|u-u_{N}\|\sim N^{-2}$ for both norms.
@@
 Method M1 and M2 produce the same values (to four points of accuracy).
-The computations with the M2 method were terminated after 10 0001000010\,00010 000 iterations, and the computations with M1 was terminated after 300 000300000300\,000300 000 iterations, marked by * in the table.
+The computations with the M2 method were terminated after 10 00010\,000 iterations, and the computations with M1 was terminated after 300 000300\,000 iterations, marked by * in the table.
 The results for the wide stencil finite difference discretizations (4th–6th column) are taken from Oberman [27].
-$N$ Bellman M1, M2 9pt stencil 17pt stencil 33pt stencil
+NN Bellman M1, M2 9pt stencil 17pt stencil 33pt stencil
 21 0.2917 0.2892 0.3115 0.2815 0.2839
@@
 511 0.2628 *
-Table 3.
-Running time for the constant Monge–Ampère example.
-Computations marked by * were terminated after 10 0001000010\,00010 000 (for method M2) and 300 000300000300\,000300 000 (for M1) iterations.
+Table 3. Running time for the constant Monge–Ampère example. Computations marked by * were terminated after 10 00010\,000 (for method M2) and 300 000300\,000 (for M1) iterations.
 Bellman M1 M2
-$N$ iter time (s) iter time (s) iter time (s)
+NN iter time (s) iter time (s) iter time (s)
 $31$ $17$ $0.1$ $2902$ $4.4$ $438$ $6.7$
-$41$ $19$ $0.5$ $5086$ $14.3$ $1539$ $23.2$
-$61$ $19$ $1.6$ 11 1461114611\,14611 146 $69.9$ $3224$ $150.3$
-$81$ $21$ $3.8$ 19 3791937919\,37919 379 $222.5$ $5443$ $546.8$
-$101$ $19$ $5.9$ 29 6982969829\,69829 698 $518.3$ $8158$ $1513.7$
-$127$ $20$ $10.8$ 46 1524615246\,15246 152 $1297.7$
-$255$ $29$ $102.7$ 174 401174401174\,401174 401 20 896.020896.020\,896.020 896.0 * *
-$511$ $25$ $870.7$
+$41$ 1919 $0.5$ $5086$ $14.3$ $1539$ $23.2$
+$61$ 1919 $1.6$ 11 14611\,146 $69.9$ $3224$ $150.3$
+$81$ 2121 $3.8$ 19 37919\,379 $222.5$ $5443$ $546.8$
+$101$ 1919 $5.9$ 29 69829\,698 $518.3$ $8158$ $1513.7$
+$127$ 2020 $10.8$ 46 15246\,152 $1297.7$ * *
+$255$ $29$ $102.7$ 174 401174\,401 20 896.020\,896.0 * *
+$511$ $25$ $870.7$ * * * *
 Running the Bellman algorithm on Equation (6.2) produces a function that (barely) fails to be convex near the corners of $\Omega$, but the same is true for Method M2, see Figure 6.
@@
 The wide stencils methods from Oberman [27] are known to converge monotonically to a viscocity supersolution of Equation (2.1), and are thus too large.
-If the solution produced by the Bellman method had been convex, Proposition 3.3 would have implied that it too is too large, but since convexity fails near the corners of the square, we have no guarantee for this.
+If the solution produced by the Bellman method had been convex, monotonicity would have implied that it too is too large, but since convexity fails near the corners of the square, we have no guarantee for this.
 Our method produces a solution which is slightly larger than the M2 and M1 methods, see Table 2.
@@
 Already on a $101\times 101$ grid, the Bellman method runs more than 300 times faster than the M2 method and about 80 times faster than M1.
-6.3. Singular examples
-6.3.1. Circular degeneracy
 (a) Error in supremum and $\ell^{2}$ norm.
@@
 $u(x,y)=\frac{1}{2}(\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}-0.2)^{+}$.
+6.3. Singular examples
+6.3.1. Circular degeneracy
 We consider
@@
 Curiously enough, the Bellman method seems to enjoy a relative speed-up as well as an improvement in accuracy for larger meshes, see Figure 8.
-If we instead do the computations on a slightly smaller square, e.g. $[0,0.99]^{2}$, the three methods run at full accuracy ($\sim N^{-2}$), and the Bellman method runs 2–3 times faster than M2.
+If we instead do the computations on a slightly smaller square, e.g. $0,0.99]^{2}$, the three methods run at full accuracy ($\sim N^{-2}$), and the Bellman method runs 2–3 times faster than M2.
 (a) Error in supremum (top) and $\ell^{2}$ norm (bottom).
@@
 7. Conclusions
-[…]
+The Bellman algorithm relies heavily on the assumption that the solution to the Monge–Ampère equation is strictly convex on the domain.
+When this assumption is satisfied, the performance of the algorithm is very good.
+Using an $N\times N$ mesh, the discretization error in supremum and $\ell^{2}$ norm is proportional to $N^{-2}$, and is achieved with few iterations (typically fewer than $10$) resulting in a running time which is proportional to $N^{2.7}$$N^{2.8}$.
 Similar performance was also observed in cases where the right hand side is mildly degenerate, such as the set where $f=0$ is at most one dimensional.